# Nuclear Astrophysics

### v1

@@ 4.1 Nucleosynthesis events and their ejecta
Stellar photospheric spectroscopy in the optical window provides for more-easily accessible abundance measurements, such as the absorption line spectroscopy of sunlight that led to the results shown in Figure  9; this has been the classical backbone of cosmic abundance data. The underlying star (here our Sun) provides the background light source that illuminates photospheric gas from below. In the early years of the 19th century, Joseph Fraunhofer and John Wollaston had noticed and studied striking absorption features in optical spectra (about 400-700 nm). With Robert Bunsen and Gustav Kirchhoff’s work about thermal radiation and its properties in heated solids, and Wilhelm Hittorf’s and Julius Plücker’s application to gas in discharge devices, this led to the spectroscope instrument (1859), a tool for then consolidating the interpretation of absorption lines representing abundances of specific elements in the solar photosphere. Since then, starlight has been used to infer elemental abundances across the table of elements, using telescopes across all countries and in space. In contrast to the nuclear gamma-ray lines, spectral line signatures at lower energies arise from electronic transitions in atomic or molecular shells. Gas temperature and ionization states determine the line features, and need to be accounted for in conversion into elemental or molecular abundances. Spectral resolution has improved substantially from technological advances, and recently could resolve the minor offsets in spectral lines that occur for different isotopes of a few elements in exceptionally-favorable cases Limitations occur from systematic effects, as the underlying radiation and the state of absorbing gas must be modeled for proper interpretation of line features. The existence or deviations from local thermodynamic equilibrium plays a role, then motions of gas in the photosphere, and the atmospheric pressure needs to be modeled properly. Opacities are complex due to the many atomic and molecular transitions that may be involved. The radiation transport through a stellar atmosphere is a complex problem at the microscopic level, and solutions used in stellar spectroscopy depend on a mix of laboratory spectroscopy, theory including ab-initio quantum-mechanical calculations, and empirical prescriptions [e.g. 35, for a recent discussion]. Classical methods such as equivalent-width measurements are often too imprecise and depend on questionable assumptions [see 41, 102, 48, for a discussion of such limitations and current stellar modeling with corresponding improvements]. A wide range of techniques have been developed for the connections to stellar parameters, formulating complex models to iteratively derive abundances [e.g. 354, 467, 137], also including machine learning [e.g. 54] as well as hierarchy concepts [234].
@@
Despite these caveats, elemental photospheric abundances have been successfully used to constrain cosmic nuclear reactions: (i) for elements that only have radioactive isotopes; the prominent example is Tc, the abundance of which was measured in AGB stars as a first proof of currently-ongoing nucleosynthesis [331]; or, as another example, (ii) for elements which have isotopes that preferentially can be attributed to either the r-process or the s process. Exploiting magic number peaks of nucleosynthesis in the s-process [see discussion in review by 311], 88Sr, 89Y, 90Zr (first s-process abundance peak), with 82, 100, and 51% of fractional isotope abundance for each element, can usefully be assigned to s process. Similarly, 139La and 140Ce (99.9 and 88%, respectively) (2nd peak), and 208Pb (62%) (3rd peak) are key isotopes where elemental abundances reflect s-process origin. Correspondingly, abundance peaks for Se, Xe, and Pt result after $\beta$ decays modify production of magic-number nuclei by the r process and its high neutron flux exposures, although always significant s-process contributions also are to be expected in these cases. Beyond the peaks resulting from magic number nuclei, elements may have specific isotopes that are shielded from the $\beta$-decay path (s-only), or separated from the valley of stability by one or more unstable isotopes along the neutron capture path (r-only), with significant relative abundance within the element to allow some conclusions on the respective process. Prominent cases are 96,94Zr (17%) versus 90-92Zr, or 76Ge (7%) versus 70-74Ge. As lifetimes of unstable nuclei above Pb and Bi are short compared to s-process time scales, all elements above Pb and Bi should have r process origins exclusively, and can be attributed unambiguously.
@@
Abundance signals from absorption lines can be obtained similarly with other background light sources. Of particular importance are background stars beyond interstellar gas clouds, or quasars or gamma-ray bursts as background sources illuminating gas from behind. At redshifts beyond z=2 the hydrogen Lyman-$\alpha$ line (121.6 nm at rest) becomes accessible to optical spectroscopy. Spectra from high-redshift quasars and $\gamma$-ray bursts with absorption lines shortward of the Lyman $\alpha$ line from the source may come from absorbing gas clouds on the near side of these sources, allowing a study of gas composition in the early universe [see 372, for a review]. In particular for determination of primordial abundances and their interpretations with respect to big-bang nucleosynthesis, this has been important for measuring helium isotopes and deuterium in gas clouds against high-redshift quasar background light [380, 381].
@@ 4.2 Sampling cosmic matter
Stardust also may directly enter the Earth atmosphere and reach surface. As in the case of meteorites, it is only under special circumstances that we can identify stardust deposits on Earth or Moon. Moreover, the transport of interstellar dust is mediated by the magnetic fields of Sun and Earth, because interstellar dust grains are charged; it is difficult to calculate the magnitude of such shielding and transport efficiency [143, 157]. A main idea was to search for some kinds of un-natural radioactivity, such as from the longer-lived actinides 244Pu (80.6 My) and 247Cm (15.7 My) [144]. A search in ice cores and deep-sea sediments for deposited dust grains from nearby nucleosynthesis events was motivated as a side activity to measure in this way the history of increased production of cosmogenic nuclides (e.g. 10Be) in the Earth atmosphere, as expected from cosmic ray intensity variations. Target isotopes were 10Be, 26Al, 36Cl, 53Mn, 60Fe, and 59Ni, and also the longer-lived 129I, 146Sm, and 244Pu [130], and the way to find very small admixtures in sediments was Accelerator Mass Spectrometry, reaching a 10-17 fraction of contaminations [see 274, for a review of the AMS technique]. Analyses of samples of ocean crust obtained from deep ocean explorations far from contaminations of civilization, and also from Antarctic snow as well as lunar samples, have been investigated [254, 147, 492, 493, 308, 141, 263, 259, 494]. The detections of 60Fe can be dated using a combination of 10Be deposits and other dating methods, to display a deposition history shown in Figure 15. Evidently, in the period around 2–3 My ago there is an enhanced deposition [493, 142]. The temporal width of this deposition appears incompatible with a debris wave from a single supernova event, and it is debated if this signature may be from multiple supernovae or from relative motion between the boundary of the Local Bubble cavity and the solar system. In any case, this graph demonstrates that deposition of rather fresh ejecta material from nearby nucleosynthesis activity occurs, and can be measured. Furthermore, a search for 244Pu ($\tau=$80 My) also recently has been successful [492, 494, 142], and indicates a correlation with the 3 My deposition peak; a rather constant influx of this radioactive species, relative to the peak-like signature from 60Fe, can however not be excluded, considering uncertainties [494]. Possible interpretations have been discussed using simulations of interstellar dynamics [428]. An origin of this 60Fe deposition from stars and supernovae of the Scorpius-Centaurus stellar group appears plausible.
@@ 4.3 Cumulative nucleosynthesis
Millions of stellar spectra have been obtained from spectroscopic surveys (SDSS/APOGEE [4], GALAH [55], Gaia-ESO [177], LAMOST [312, 535], and RAVE [451]) with multi-fiber medium and high-resolution spectrographs [e.g. 121, 264]. Analysis often is conducted through automatic pipelines, using empirical abundance fitting engines such as ’The Cannon’ [63, 353] and benchmark stars for calibrations and then deriving abundance results. From these, it has been concluded that the Galaxy’s inner region and bulge, as well as the halo, is mostly composed of old stars that were created in the early phase of the Galaxy’s evolution, while the disk and spiral-arm regions include stars of all ages, and show the result of ongoing star formation over the Galactic history.
A description of such compositional evolution has been developed from the early work of Tinsley [468]. This work expanded into various models of Cosmic (or Galactic) Chemical [^7] We note that ’chemical’ here refers to the chemical elements; no chemical reactions are addressed in CCE/GCE models. Evolution describing the increase of metal abundances in terms of the star formation rate and the yields of the variety of nucleosynthesis sources [81, 386, 176, 324, 257, 325]. The sources that are included here are mainly stars and their supernovae of all types I and II, but also rarer sources from binary systems such as neutron-star mergers, as well as rare supernova species such as magnetic-jet supernovae and hypernovae, have been included. Herein, the delay times between star formation and the ejection of fresh nucleosynthesis material from the respective sources provide an important ingredient, and rely on astrophysical theory or modeling. Figure 16 shows the performance of a one-zone model for the solar neighborhood, with infall at a characteristic time scale of 10 Gy, comparing it to a large assembly of photospheric abundance data [for details see 388].
@@
Chemical-evolution models have advanced from one-zone models to 2- and 3-dimensional. Although not based on fundamental-physics modelings, these descriptions provide a framework, in which the different components of nuclear astrophysics can be linked together to provide a bridge to astronomical observations [see reviews by 325, 256, 110]. Analytical descriptions are complemented by numerical approaches such as smooth-particle hydrodynamics, which enable also inhomogeneous chemical evolution prescriptions. In this may, links appear to simulations of galactic evolution, and stimulate cross comparison between the evolution of galactic structures and of compositional evolution. The determination of stellar ages [see 445, for a recent review] provides one of the major uncertainties in the data themselves, comparisons across selection biases are important. Since the precision kinematic measurements of Hipparcos [101] and in particular the Gaia mission [162], abundances can be compared with kinematic data, and a chemical tagging of kinematic groups obtains new structural information [e.g. 61, 127]. This revealed that the Galaxy’s history included collisions with smaller galaxies, disturbing the gradual build-up of abundances beyond primordial composition from nucleosynthesis in stars and their explosions, thus confirming earlier detections of stellar streams [202].
@@
A completely independent way to study the cumulative effects of nucleosynthesis is the field of cosmochemistry [537, 82, 357]. Here, material samples are analyzed and interpreted in terms of cumulative nucleosynthesis. The inclusions of pre-solar stardust grains and their interpretation with respect to specific sources of nucleosynthesis has been discussed above (section 4.2); this was based on the plausible assumption of such stardust identified by strikingly-different isotopic signatures having condensed in the vicinity of such a sources, thus freezing-in and preserving its characteristic ejecta [see, however, 355, for cosmic evolution effects in stardust]. The sample of interstellar gas that was isolated as the solar system formed is preserved by all matter attributed to the solar system, including meteorites (their bulk material, beyond the specific stardust inclusions); it carries the complete nucleosynthetic history of the past, mediated by material flows connected to it. As described above, the solar abundances shown in Figure 9 are a composite result of spectroscopy of solar radiation and of mass spectrometry of meteoritic material, the latter essentially contribution all isotopic abundance information [305, 306]. On Earth, abundances of elements and isotopes have been altered through a combination of physical, chemical, and antropomorphic processes; the abundances in the crust of our planet deviate significantly from the distribution shown in Figure 9, with depletion of gases and enrichments of refractory elements. The microscopic fractionation processes that may have altered meteorite compositions over their history of 4.6 Gy are complex, but mostly rather well understood [92]. The interactions and dynamics caused by the planets within the solar system lead to macroscopic effects of some second-order heterogeneity across the meteoritic samples [253]. Nevertheless, meteorites are a valuable repository for cosmochemistry, as discussed in the review by Nittler and Dauphas [357]. As in the case of stardust, the key is that mass spectrometry in terrestrial laboratories provides access to isotopic information for the variety of even rarest elements.
@@ 4.4 Constraining stellar matter in compact stars
Oscillation frequencies measured at stellar surfaces provide tools to indirectly constrain the interiors of a star [see 9, for a recent review of Astero-seismology]. In the case of the Sun, one of the parameters driving stellar oscillations is the depth of the convective zone, which sensitively depends on the composition of matter, specifically their metallicity, or abundances of main elements Carbon, Nitrogen, and Oxygen [26, 488]. Helioseismology has obtained a precision of 0.5% and thus enabled detailed constraints on the depth of the convection zone in the Sun. This can be compared to solar model predictions, evaluating uncertainties in solar composition, and how this depends on nuclear reaction rates [6]. Various revisions of the solar metallicity since the value of 0.0187 from the early standard reference [16] have been made [25, 26, 27], and currently converge on a metallicity value of 0.0177 [314], alleviating the earlier reports of a major reduction to values near 0.014, and now consistent with helioseismic constraints. With the Kepler and TESS missions, astero-seismology has become a tool to better constrain stellar ages [106], as internal structure changes with consumption of fuel in nuclear burning. Similarly, the pulsations of a white dwarf star reflects the internal composition. In this way, specifically the central Oxygen content, could be derived [174]. Thus astero-seismology had an important impact on studies of Galactic archeology. Improving age estimates for stellar groups [448, 449, 439], this is a key step to disentangle stellar groups of different origins, as the degeneracy between age effects and different star forming histories is lifted.

### v2

@@ 4.1 Nucleosynthesis events and their ejecta
Stellar photospheric spectroscopy in the optical window provides for more-easily accessible abundance measurements, such as the absorption line spectroscopy of sunlight that led to the results shown in Figure  9; this has been the classical backbone of cosmic abundance data. The underlying star (here our Sun) provides the background light source that illuminates photospheric gas from below. In the early years of the 19th century, Joseph Fraunhofer and John Wollaston had noticed and studied striking absorption features in optical spectra (about 400-700 nm). With Robert Bunsen and Gustav Kirchhoff’s work about thermal radiation and its properties in heated solids, and Wilhelm Hittorf’s and Julius Plücker’s application to gas in discharge devices, this led to the spectroscope instrument (1859), a tool for then consolidating the interpretation of absorption lines representing abundances of specific elements in the solar photosphere. Since then, starlight has been used to infer elemental abundances across the table of elements, using telescopes across all countries and in space. In contrast to the nuclear gamma-ray lines, spectral line signatures at lower energies arise from electronic transitions in atomic or molecular shells. Gas temperature and ionization states determine the line features, and need to be accounted for in conversion into elemental or molecular abundances. Spectral resolution has improved substantially from technological advances, and recently could resolve the minor offsets in spectral lines that occur for different isotopes of a few elements in exceptionally-favorable cases Limitations occur from systematic effects, as the underlying radiation and the state of absorbing gas must be modeled for proper interpretation of line features. The existence or deviations from local thermodynamic equilibrium plays a role, then motions of gas in the photosphere, and the atmospheric pressure needs to be modeled properly. Opacities are complex due to the many atomic and molecular transitions that may be involved. The radiation transport through a stellar atmosphere is a complex problem at the microscopic level, and solutions used in stellar spectroscopy depend on a mix of laboratory spectroscopy, theory including ab-initio quantum-mechanical calculations, and empirical prescriptions [e.g. 36, for a recent discussion]. Classical methods such as equivalent-width measurements are often too imprecise and depend on questionable assumptions [see 42, 103, 49, for a discussion of such limitations and current stellar modeling with corresponding improvements]. A wide range of techniques have been developed for the connections to stellar parameters, formulating complex models to iteratively derive abundances [e.g. 357, 471, 138], also including machine learning [e.g. 55] as well as hierarchy concepts [235].
@@
Despite these caveats, elemental photospheric abundances have been successfully used to constrain cosmic nuclear reactions: (i) for elements that only have radioactive isotopes; the prominent example is Tc, the abundance of which was measured in AGB stars as a first proof of currently-ongoing nucleosynthesis [333]; or, as another example, (ii) for elements which have isotopes that preferentially can be attributed to either the r-process or the s process. Exploiting magic number peaks of nucleosynthesis in the s-process [see discussion in review by 313], 88Sr, 89Y, 90Zr (first s-process abundance peak), with 82, 100, and 51% of fractional isotope abundance for each element, can usefully be assigned to s process. Similarly, 139La and 140Ce (99.9 and 88%, respectively) (2nd peak), and 208Pb (62%) (3rd peak) are key isotopes where elemental abundances reflect s-process origin. Correspondingly, abundance peaks for Se, Xe, and Pt result after $\beta$ decays modify production of magic-number nuclei by the r process and its high neutron flux exposures, although always significant s-process contributions also are to be expected in these cases. Beyond the peaks resulting from magic number nuclei, elements may have specific isotopes that are shielded from the $\beta$-decay path (s-only), or separated from the valley of stability by one or more unstable isotopes along the neutron capture path (r-only), with significant relative abundance within the element to allow some conclusions on the respective process. Prominent cases are 96,94Zr (17%) versus 90-92Zr, or 76Ge (7%) versus 70-74Ge. As lifetimes of unstable nuclei above Pb and Bi are short compared to s-process time scales, all elements above Pb and Bi should have r process origins exclusively, and can be attributed unambiguously.
@@
Abundance signals from absorption lines can be obtained similarly with other background light sources. Of particular importance are background stars beyond interstellar gas clouds, or quasars or gamma-ray bursts as background sources illuminating gas from behind. At redshifts beyond z=2 the hydrogen Lyman-$\alpha$ line (121.6 nm at rest) becomes accessible to optical spectroscopy. Spectra from high-redshift quasars and $\gamma$-ray bursts with absorption lines shortward of the Lyman $\alpha$ line from the source may come from absorbing gas clouds on the near side of these sources, allowing a study of gas composition in the early universe [see 375, for a review]. In particular for determination of primordial abundances and their interpretations with respect to big-bang nucleosynthesis, this has been important for measuring helium isotopes and deuterium in gas clouds against high-redshift quasar background light [383, 384].
@@ 4.2 Sampling cosmic matter
Stardust also may directly enter the Earth atmosphere and reach surface. As in the case of meteorites, it is only under special circumstances that we can identify stardust deposits on Earth or Moon. Moreover, the transport of interstellar dust is mediated by the magnetic fields of Sun and Earth, because interstellar dust grains are charged; it is difficult to calculate the magnitude of such shielding and transport efficiency [144, 158]. A main idea was to search for some kinds of un-natural radioactivity, such as from the longer-lived actinides 244Pu (80.6 My) and 247Cm (15.7 My) [145]. A search in ice cores and deep-sea sediments for deposited dust grains from nearby nucleosynthesis events was motivated as a side activity to measure in this way the history of increased production of cosmogenic nuclides (e.g. 10Be) in the Earth atmosphere, as expected from cosmic ray intensity variations. Target isotopes were 10Be, 26Al, 36Cl, 53Mn, 60Fe, and 59Ni, and also the longer-lived 129I, 146Sm, and 244Pu [131], and the way to find very small admixtures in sediments was Accelerator Mass Spectrometry, reaching a 10-17 fraction of contaminations [see 275, for a review of the AMS technique]. Analyses of samples of ocean crust obtained from deep ocean explorations far from contaminations of civilization, and also from Antarctic snow as well as lunar samples, have been investigated [255, 148, 496, 497, 310, 142, 264, 260, 498]. The detections of 60Fe can be dated using a combination of 10Be deposits and other dating methods, to display a deposition history shown in Figure 15. Evidently, in the period around 2–3 My ago there is an enhanced deposition [497, 143]. The temporal width of this deposition appears incompatible with a debris wave from a single supernova event, and it is debated if this signature may be from multiple supernovae or from relative motion between the boundary of the Local Bubble cavity and the solar system. In any case, this graph demonstrates that deposition of rather fresh ejecta material from nearby nucleosynthesis activity occurs, and can be measured. Furthermore, a search for 244Pu ($\tau=$80 My) also recently has been successful [496, 498, 143], and indicates a correlation with the 3 My deposition peak; a rather constant influx of this radioactive species, relative to the peak-like signature from 60Fe, can however not be excluded, considering uncertainties [498]. Possible interpretations have been discussed using simulations of interstellar dynamics [431]. An origin of this 60Fe deposition from stars and supernovae of the Scorpius-Centaurus stellar group appears plausible.
@@ 4.3 Cumulative nucleosynthesis
Millions of stellar spectra have been obtained from spectroscopic surveys (SDSS/APOGEE [5], GALAH [56], Gaia-ESO [178], LAMOST [314, 539], and RAVE [454]) with multi-fiber medium and high-resolution spectrographs [e.g. 122, 265]. Analysis often is conducted through automatic pipelines, using empirical abundance fitting engines such as ’The Cannon’ [64, 356] and benchmark stars for calibrations and then deriving abundance results. From these, it has been concluded that the Galaxy’s inner region and bulge, as well as the halo, is mostly composed of old stars that were created in the early phase of the Galaxy’s evolution, while the disk and spiral-arm regions include stars of all ages, and show the result of ongoing star formation over the Galactic history.
A description of such compositional evolution has been developed from the early work of Tinsley [472]. This work expanded into various models of Cosmic (or Galactic) Chemical [^7] We note that ’chemical’ here refers to the chemical elements; no chemical reactions are addressed in CCE/GCE models. Evolution describing the increase of metal abundances in terms of the star formation rate and the yields of the variety of nucleosynthesis sources [82, 389, 177, 326, 258, 327]. The sources that are included here are mainly stars and their supernovae of all types I and II, but also rarer sources from binary systems such as neutron-star mergers, as well as rare supernova species such as magnetic-jet supernovae and hypernovae, have been included. Herein, the delay times between star formation and the ejection of fresh nucleosynthesis material from the respective sources provide an important ingredient, and rely on astrophysical theory or modeling. Figure 16 shows the performance of a one-zone model for the solar neighborhood, with infall at a characteristic time scale of 10 Gy, comparing it to a large assembly of photospheric abundance data [for details see 391].
@@
Chemical-evolution models have advanced from one-zone models to 2- and 3-dimensional. Although not based on fundamental-physics modelings, these descriptions provide a framework, in which the different components of nuclear astrophysics can be linked together to provide a bridge to astronomical observations [see reviews by 327, 257, 111]. Analytical descriptions are complemented by numerical approaches such as smooth-particle hydrodynamics, which enable also inhomogeneous chemical evolution prescriptions. In this may, links appear to simulations of galactic evolution, and stimulate cross comparison between the evolution of galactic structures and of compositional evolution. The determination of stellar ages [see 448, for a recent review] provides one of the major uncertainties in the data themselves, comparisons across selection biases are important. Since the precision kinematic measurements of Hipparcos [102] and in particular the Gaia mission [163], abundances can be compared with kinematic data, and a chemical tagging of kinematic groups obtains new structural information [e.g. 62, 128]. This revealed that the Galaxy’s history included collisions with smaller galaxies, disturbing the gradual build-up of abundances beyond primordial composition from nucleosynthesis in stars and their explosions, thus confirming earlier detections of stellar streams [203].
@@
A completely independent way to study the cumulative effects of nucleosynthesis is the field of cosmochemistry [541, 83, 360]. Here, material samples are analyzed and interpreted in terms of cumulative nucleosynthesis. The inclusions of pre-solar stardust grains and their interpretation with respect to specific sources of nucleosynthesis has been discussed above (section 4.2); this was based on the plausible assumption of such stardust identified by strikingly-different isotopic signatures having condensed in the vicinity of such a sources, thus freezing-in and preserving its characteristic ejecta [see, however, 358, for cosmic evolution effects in stardust]. The sample of interstellar gas that was isolated as the solar system formed is preserved by all matter attributed to the solar system, including meteorites (their bulk material, beyond the specific stardust inclusions); it carries the complete nucleosynthetic history of the past, mediated by material flows connected to it. As described above, the solar abundances shown in Figure 9 are a composite result of spectroscopy of solar radiation and of mass spectrometry of meteoritic material, the latter essentially contribution all isotopic abundance information [306, 307]. On Earth, abundances of elements and isotopes have been altered through a combination of physical, chemical, and antropomorphic processes; the abundances in the crust of our planet deviate significantly from the distribution shown in Figure 9, with depletion of gases and enrichments of refractory elements. The microscopic fractionation processes that may have altered meteorite compositions over their history of 4.6 Gy are complex, but mostly rather well understood [93]. The interactions and dynamics caused by the planets within the solar system lead to macroscopic effects of some second-order heterogeneity across the meteoritic samples [254]. Nevertheless, meteorites are a valuable repository for cosmochemistry, as discussed in the review by Nittler and Dauphas [360]. As in the case of stardust, the key is that mass spectrometry in terrestrial laboratories provides access to isotopic information for the variety of even rarest elements.
@@ 4.4 Constraining stellar matter in compact stars
Oscillation frequencies measured at stellar surfaces provide tools to indirectly constrain the interiors of a star [see 10, for a recent review of Astero-seismology]. In the case of the Sun, one of the parameters driving stellar oscillations is the depth of the convective zone, which sensitively depends on the composition of matter, specifically their metallicity, or abundances of main elements Carbon, Nitrogen, and Oxygen [27, 492]. Helioseismology has obtained a precision of 0.5% and thus enabled detailed constraints on the depth of the convection zone in the Sun. This can be compared to solar model predictions, evaluating uncertainties in solar composition, and how this depends on nuclear reaction rates [7]. Various revisions of the solar metallicity since the value of 0.0187 from the early standard reference [17] have been made [26, 27, 28], and currently converge on a metallicity value of 0.0177 [316], alleviating the earlier reports of a major reduction to values near 0.014, and now consistent with helioseismic constraints. With the Kepler and TESS missions, astero-seismology has become a tool to better constrain stellar ages [107], as internal structure changes with consumption of fuel in nuclear burning. Similarly, the pulsations of a white dwarf star reflects the internal composition. In this way, specifically the central Oxygen content, could be derived [175]. Thus astero-seismology had an important impact on studies of Galactic archeology. Improving age estimates for stellar groups [451, 452, 442], this is a key step to disentangle stellar groups of different origins, as the degeneracy between age effects and different star forming histories is lifted.